### v2

# ProjectTest: A Project-level LLM Unit Test Generation Benchmark and Impact of Error Fixing Mechanisms

Abstract
Unit test generation has become a promising and important use case of LLMs. However, existing evaluation benchmarks for assessing LLM unit test generation capabilities focus on function- or class-level code rather than more practical and challenging project-level codebases. To address such limitation, we propose ProjectTest, a project-level benchmark for unit test generation covering Python, Java, and JavaScript. ProjectTest features 20 moderate-sized and high-quality projects per language. We evaluate nine frontier LLMs on ProjectTest and the results show that all frontier LLMs tested exhibit moderate performance on ProjectTest on Python and Java, highlighting the difficulty of ProjectTest. We also conduct a thorough error analysis, which shows that even frontier LLMs, such as Claude-3.5-Sonnet, have significant simple errors, including compilation and cascade errors. Motivated by this observation, we further evaluate all frontier LLMs under manual error-fixing and self-error-fixing scenarios to assess their potential when equipped with error-fixing mechanisms.
[…]
ProjectTest: A Project-level LLM Unit Test Generation Benchmark and Impact of Error Fixing Mechanisms
Yibo Wang1  Congying Xia2  Wenting Zhao3  Jiangshu Du1thanks: Work Done Prior to Amazon Chunyu Miao1Zhongfen Deng1Philip S. Yu1Chen Xing4 1University of Illinois Chicago, 3 Salesforce Research, 4 Scale AI {ywang633, jdu25, cmiao8, zdeng21, psyu}@uic.edu congyingxia3@gmail.com, wenting.zhao@salesforce.com, chen.xing@scale.com
1 Introduction
Unit testing plays an important role in software development, helping identify bugs and ensuring codes are solid and maintainable. Writing unit tests is time-consuming, usually accounting for approximately 15.8% of software development time for developers Daka and Fraser (2014). Therefore, automated test case generation, like search-based Fraser and Arcuri (2011); Harman and McMinn (2009), constraint-based (Xiao et al., 2013), and random-based (Pacheco et al., 2007) methods, has been proposed to create unit tests. However, the generated unit tests are usually less readable than manually-written tests and limited to certain types of functions Grano et al. (2018). Lately, large language models (LLMs) have become game-changers, significantly accelerating unit test generation and improving readability and generalizability with little to no human effort Siddiq et al. (2024); Xie et al. (2023).
Given the rapid adoption of LLMs on unit testing, the evaluation of LLM unit test generation capabilities appears to be lagging behind. Previous unit test generation evaluation benchmarks primarily focus on function-level, class-level, or file-level Chen et al. (2021); Du et al. (2023); Wang et al. (2024); Jain et al. (2024a) codes. However, project-level codes are more representative of real-world scenarios and practical needs. The complex dependency relationships between different files in project-level codebases make unit test generation more challenging. The only existing benchmark that has briefly explored project-level unit test generation is DevBench Li et al. (2024). However, due to its broad focus, the number of projects included for unit test generation is low for each language (e.g., 5 for Java and 5 combined for C and C++), with varying quality. Half of its projects for unit test generation evaluation are difficulty to track, and most of the identifiable projects have fewer than 250 Stars and fewer than 50 Forks. DevBench also didn’t provide a thorough analysis of error types, potentials, or self-fixing capabilities of frontier LLMs’ project-level unit test generation.
Therefore, we propose a new project-level unit test generation evaluation benchmark, ProjectTest, to offer a larger, higher-quality project set for project-level unit test generation along with a more thorough error analysis of frontier LLMs on unit test generation. ProjectTest covers three programming languages: Python, Java, and JavaScript. For each programming language, we construct 20 projects filtered from GitHub [^1] https://github.com/. ProjectTest applies clear filtering criteria to select projects. It includes moderate-sized projects with multiple files and dependencies between them. Each project has less than 1,600 lines of code, which fits within the maximum input length of most code language models. Quality is ensured by the number of stars and forks.
We evaluate nine frontier LLMs, such as Claude-3.5-Sonnet Anthropic (2024), Gemini-1.5-Pro Team et al. (2024b), and GPT-o1, on ProjectTest and conduct comprehensive error analyses. We find that all tested frontier LLMs perform moderately on ProjectTest on Python and Java, highlighting the difficulty of ProjectTest. We also observe that different LLMs have different language-level expertise. Claude-3.4-Sonnet ranks first in Java, while GPT-o1 ranks first in JavaScript. Among three programming languages, Java is the most difficult language, primarily due to stricter syntax. Among all the tested models, GPT-o1 performs the best in general, especially in JavaScript.
Error analyses from above also show that even frontier LLMs, like Claude-3.5-Sonnet, have significant compilation and cascade errors. Although these errors appear to be preliminary and may be relatively easy to fix, they prevent us from observing more advanced aspects of LLM performance on unit test generation, such as correctness and coverage. To address this, we first manually fix LLM’s compilation and cascade errors and then re-evaluate the fixed unit tests. This allows us to measure not only the models’ raw performance but also their potential for improvement when combined with simple error-fixing mechanisms. By incorporating error-fixing, we uncover critical insights into the effort required to refine generated tests and better understand the various types of errors that occur in unit tests generated by different LLMs. We observe that the model rankings change significantly after the manual fix, showing the significant differences in different LLMs’ error distribution and their potentials after error-fixing. Inspired by such findings from manual fixes, we also explore using LLMs for self-fixing their errors in generating project-level unit tests. The results show that while LLMs can correct some errors in their generated unit tests, their self-fixing abilities still lag behind the quality and reliability of human fixes.
We summarize the contribution of this paper as follows: we introduce the first project-level evaluation benchmark for unit test generation and conduct an extensive evaluation of nine frontier LLMs. Additionally, we conduct thorough error analyses by manually fixing compilation and cascade errors and provide critical insights. Inspired by the error analysis, we are the first to assess LLMs’ self-fixing capability on unit test generation.
2 Related Work
2.1 Traditional Unit Test Generation
Traditional unit test generation methods employ search-based (Harman and McMinn, 2009; Fraser and Arcuri, 2011; Lukasczyk and Fraser, 2022), constraint-based (Xiao et al., 2013), or random-based (Pacheco et al., 2007) strategies to construct test suites that maximize code coverage. Although these traditional approaches can generate unit tests with reasonable coverage, the resulting tests often have lower readability and less meaningfulness compared to developer-written tests. As a result, automatically generated tests are frequently not directly adopted by practitioners in real-world scenarios (Almasi et al., 2017; Grano et al., 2019).
2.2 LLM-enhanced Unit Test Generation
Large Language Models have demonstrated strong code generation capabilities, inspiring their use in automated unit test generation. Recent approaches in LLM-enhanced unit test generation leverage zero-shot strategies (Siddiq et al., 2024), iterative querying (Schäfer et al., 2023), fine-tuning on specialized datasets (Alagarsamy et al., 2024), adaptive context selection (Xie et al., 2023), and focusing on subtle code differences (Dakhel et al., 2024; Li et al., 2023). These methods are evaluated with various metrics—including compilation success, test correctness, coverage, and bug detection—and demonstrate that LLMs can effectively surpass traditional test generation techniques.
2.3 Unit Test Generation Benchmark
Current benchmarks for LLM-based unit test generation mainly focus on function-level Wang et al. (2024), class-level Du et al. (2023), or file-level code Jain et al. (2024a). Project-level software testing benchmarks, on the other hand, often target tasks other than unit test generation. For instance, R2E-Eval1 Jain et al. (2024b) is designed for the generation of equivalent test harnesses, SWT-Bench Mündler et al. (2024) focuses on fixing specific bugs rather than entire projects, and DevBench Li et al. (2024) centers on software development tasks. While DevBench skims the surface of project-level unit testing, its data size and quality are quite mixed, particularly for C/C# and Java, with only five projects, respectively. Additionally, given its broad focus, DevBench doesn’t include comprehensive evaluation and error analysis on project-level unit test generation.
3 Methodology
Figure 1: Overview of the unit test generation process.
We first introduce the dataset collection and preprocessing of creating ProjectTest (§3.1). Then, we introduce evaluation metrics (§3.2) and provide an overview of the Unit Test Generation pipeline (§3.3) that we use to evaluate LLMs on ProjectTest, including three unit test generation scenarios.
3.1 Benchmark Dataset
Dataset Collection
Our dataset is built from carefully selected project-level repositories on GitHub, which is a popular platform for hosting and collaborating on software development projects. We focus on three widely used programming languages: Python, Java, and JavaScript. We establish our selection criteria based on three key factors: 1) a reasonable size, 2) inter-file dependencies, and 3) a reliable source. Thus, we collect reliable and self-contained projects consisting of 2 to 15 files with fewer than 1600 lines of code (LOC). We limit our selection to repositories with publicly available licenses, such as the MIT license, ensuring the legality and openness of the code. To maintain the quality and reliability of the dataset, we choose projects with a high number of stars and forks, which signals community approval and widespread usage. For projects that fit all requirements above but are too big for current frontier LLMs to handle, we also extracted smaller projects from these large codebases. These smaller projects were carefully adjusted to be self-contained without relying on the original larger projects. After applying these criteria, we constructed 20 representative projects per programming language. The summary of dataset statistics is presented in Table 1, and detailed information on the dataset sources and statistics for each project can be found in Appendix A.
Pre-processing
Dataset pre-processing involves several key steps to ensure that the projects are well-structured and suitable for testing. First, we double-check whether the selected projects have syntax errors, even though they are sourced from reliable codebases. Second, for projects extracted from a larger codebase, we modify them to be self-contained by reorganizing files, adjusting domain naming conventions, and/or modifying import paths to remove dependencies on external modules. Next, to enhance the accuracy of line coverage measurements, we consolidate statements that are split across multiple lines into a single line, ensuring that the metrics are more valid. Additionally, we maintain the integrity of the original code style as much as possible, preserving the diverse coding practices across different projects. This approach allows us to test how LLMs handle various code styles in a realistic environment.
Table 1: ProjectTest data statistics. LOC represents lines of code.
Language Avg. #Files Avg. LOC Avg. #Stars Avg. #Forks
Python 6.10 654.60 5810.30 996.90
Java 4.65 282.60 3306.05 1347.65
JavaScript 4.00 558.05 17242.30 5476.45
3.2 Evaluation Metrics
We focus on three key aspects when evaluating the generated unit tests: compilation rate, correctness rate, and coverage rate. Compilation rate (ComR) measures the percentage of projects in which the generated test suites compile successfully, indicating how often LLMs produce unit test suites that can be executed without compilation errors. The compilation rate for all projects in $X$ is defined as $ComR=\frac{|X^{com}|}{|X|}$, where $X$ is the project set and $X^{com}\subset X$ denotes the subset of projects whose test suites compile successfully. Correctness rate (CR) calculates the percentage of unit tests that are correct out of all generated unit tests for each project, providing insight into the accuracy of the test generation process. On average, more than 95% of vanilla-generated unit tests compare expected and actual values, reinforcing the validity of CR as an evaluation metric. Detailed statistics see Appendex C. The correctness rate for the project $x$ is defined as $CR_{x}=\frac{|T_{x}^{cor}|}{|T_{x}|}$, where $T_{x}$ is the generated test suite and $T_{x}^{cor}\subset T_{x}$ denotes the correct unit test set for the project $x$. Coverage rate analyzes both line and branch coverage to understand how well the generated unit tests explore the code’s functionality. The coverage rate for the project $x$ is defined as $CR_{x}=\frac{covered(x)}{total(x)}$, where $covered(x)$ denotes the number of covered lines/branches in project $x$ and $total(x)$ the total number of lines/branches in project $x\in X$.
These three evaluation metrics are not independent. If a project has the generated test suite containing compilation errors, none of its unit tests can be executed successfully, leading to both the correctness rate and the coverage rate for the project being zeros. Additionally, some errors resulting in failed tests, like missing Python dependencies, can also lead to a change in coverage rate. Therefore, considering the interdependencies between the three evaluation metrics, we extend our analysis beyond the evaluation of vanilla unit tests to include manually fixing these errors. This enables a more comprehensive assessment of LLMs’ potential to generate high-quality unit tests once these errors are addressed. This assessment is conducted while maintaining the same quantity and diversity of unit tests originally generated by the LLMs. Furthermore, we extend our analysis to examine the self-fixing capabilities of LLMs.
{mdframed}
[style=dataset, backgroundcolor=white] Refer to caption
Figure 2: An example of ProjectTest.
3.3 Unit Test Generation
Figure 1 shows an overview of the unit test generation process by LLMs. Our unit test generation and evaluation aim to ensure fair and thorough assessments of unit tests generated by LLMs under different scenarios:
  • Scenario 1: Vanilla unit tests extracted from LLMs’ outputs.
  • Scenario 2: Compilable unit tests after manually fixing all compilation and cascade errors.
  • Scenario 3: Unit tests refined by LLMs self-fixing, provided with error messages and conversation history.
{mdframed}
[style=prompt, backgroundcolor=white] System Prompt: You are a coding assistant. You generate only source code.
User Prompt: {Original Codes} Please generate enough unit test cases for each Python file in the project. Ensure that the import path is correct, depending on whether the project is structured as a package. Make sure the tests can successfully compile. Make sure the tests have correct results. Try to achieve the highest coverage rate.
Figure 3: The prompt used to generate unit tests for Python projects. Purple indicates language-specific instruction. Blue, orange, and red indicates instructions related to compilation rate, correctness rate, and coverage rate, respectively.
Scenario 1: Vanilla Unit Test Generation. We begin by inputting the entire project and the carefully crafted prompt into the LLM, ensuring the context and requirements are clearly communicated. The complete project codes are used as input to ensure the models have all the necessary context to generate unit tests for the entire project. An example is shown in Figure 2. We carefully design prompts for different LLMs to thoroughly test their actual capabilities. In addition, to address specific issues associated with different programming languages, we incorporate language-specific prompts tailored to solve particular challenges. We require the LLMs to generate unit tests for each file within the project. Furthermore, we provide detailed prompts instructing LLMs to focus on various evaluation aspects, including compilation rate, correctness rate, and coverage rate. This structured prompt engineering enhances the effectiveness and relevance of the outputs produced by the LLMs. An example of our designed prompt for Python is shown in Figure 3. All prompts used in our experiments are listed in Appendix B.1, and an ablation analysis of the prompts is shown in Appendix D.1. The vanilla unit tests are extracted from the LLM response based on the input project and prompt.
Scenario 2: Manual Fixing compilation and cascade errors. Manually fixing compilation and cascade errors is motivated by our empirical observation from scenario 1 that even the vanilla unit tests from state-of-the-art LLMs, such as Claude-3.5-Sonnet, contain significant compilation errors, making them non-compilable. Additionally, they exhibit cascade errors that are easy to fix but can affect multiple unit tests or the entire test suite (details in Section 5.5). Although these errors are preliminary and relatively simple to resolve, they hinder further analysis of other aspects of LLM performance on unit test generation, such as correctness and coverage.
{mdframed}
[style=errors, backgroundcolor=white] Refer to caption
Figure 4: An example of compilation error generated by GPT-4-Turbo.
{mdframed}
[style=errors, backgroundcolor=white] Refer to caption
Figure 5: An example of cascade error generated by CodeQwen1.5-7B-Chat.
Therefore, based on vanilla unit tests, we make the minimum necessary changes to resolve compilation errors and cascade errors, focusing solely on eliminating these errors without altering the original test intent. Compilation errors are defined as errors that prevent testing frameworks from executing. [^2] Technically, Python does not require compilation. We refer to errors that cause pytest to fail before it can collect and run any tests as compilation errors. As shown in Figure 4, the ModuleNotFoundError causes pytest to fail before collecting any unit tests, making the entire test suite uncompilable. This results not only in compilation failure but also in unreachable correctness and coverage rates. [^3] We consider unreachable correctness and coverage rate as zero. Cascade errors are defined as errors that cause cascading failures across multiple unit tests or even the entire test suite. As shown in Figure 5, although the tests are fundamentally correct, this NameError (missing NumPy) invalidates multiple or even the entire test suite. By resolving these errors, manual fixing ensures that all unit tests are compilable and no cascade errors invalidate tests that are fundamentally correct.
Manually fixing compilation and cascade errors plays a crucial role in evaluating the quality and reliability of generated unit tests. By addressing these errors, we gain deeper insights into the effectiveness of LLM-generated unit tests and identify areas for improvement. This process also helps assess the potential for LLMs to improve continuously once such simple errors are resolved. Additionally, we evaluate unit tests fixing only compilation errors in Appendix D.2.
Scenario 3: LLM Self-fixing
{mdframed}
[style=prompt, backgroundcolor=white] System Prompt: You are a coding assistant…
User Prompt: {Original Codes} Please generate enough unit test cases…
LLM Response: {Generated Vanilla Unit Tests}
User Prompt: Here are the error messages from the tests: {Error Messages}. Errors exist in the generated unit tests. Please fix the unit tests to address these errors and provide the entire unit tests.
Figure 6: The prompt used for the LLM self-fixing scenario for Python projects.
Inspired by our observation from manual fixing that different LLMs exhibit significantly different potentials after manual fixing, we are curious about how LLMs would perform in self-fixing on our benchmark. We explore LLMs’ self-fixing abilities by incorporating conversation history and error messages as shown in Figure 6. We provide LLMs with the conversation history (including the system prompt, the user prompt for unit test generation requests, and LLM vanilla response), error messages obtained from the testing framework, and the user prompt for error fixing requests. When the open-source LLM’s input length is limited, we prioritize the information in the following order: system prompt, LLM’s initial response, error messages, error-fixing requests, and unit test generation requests. Less important information is truncated as needed. Additionally, we reserve at least 2,000 tokens for the open-source LLM’s self-fixing outputs. LLM self-fixing scenario helps us understand LLMs’ error-fixing ability and their potential to generate better unit tests when incorporating the self-fixing process. Note that during self-fixing, we do not constrain the target error types to just compilation or cascade errors.
4 Experimental Settings
4.1 Models
We evaluate five close-sourced models: GPT-o1, Gemini-1.5-Pro Team et al. (2024b), Claude-3.5-Sonnet-20241022 (Claude-3.5-Sonnet) Anthropic (2024), GPT-4-Turbo Achiam et al. (2023) and GPT-3.5-Turbo, and four open-sourced models: CodeQwen1.5-7B-Chat (CodeQwen1.5) Bai et al. (2023), DeepSeek-Coder-6.7b-Instruct (DeepSeek-Coder) Guo et al. (2024); Zhu et al. (2024), CodeLlama-7b-Instruct-hf (CodeLlama) Roziere et al. (2023), and CodeGemma-7b-it (CodeGemma) Team et al. (2024a). Detailed information is in Appendix B.2.
4.2 Implementation Details
We use zero-shot prompting for unit test generation. The temperature is set to 0 during inference. Experiments are conducted on 8 NVIDIA A100 GPUs. The maximum input length is configured to match the token limit of each LLM to evaluate model capabilities. We use Pytest [^4] https://docs.pytest.org/en/stable/ for Python, Jacoco [^5] https://www.eclemma.org/jacoco/ for Java, and JEST [^6] https://jestjs.io/ for JavaScript regarding testing frameworks.
5 Experiments
We evaluate the generated unit tests from three scenarios, vanilla (§ 5.1), after manual fixing of compilation and cascade errors (§ 5.2), and LLM self-fixing (§ 5.3). For each scenario, we evaluate the Correctness Rate (CR), Compilation Rate (ComR), Line Coverage (LC), and Branch Coverage (BC). We also conduct unique contribution analyses (§5.4) and detailed error analyses (§ 5.5).
5.1 Main Results
Table 2: Main Results. CR represents Correctness Rate; ComR represents Compilation Rate; LC represents Line Coverage; BC represents Branch Coverage.
Language Model CR ComR LC BC #Tests #Correct
Python GPT-4-Turbo 47% 65% 40% 36% 12.60 6.15
GPT-3.5-Turbo 37% 60% 38% 34% 16.90 6.65
GPT-o1 60% 65% 56% 54% 36.35 21.7
Gemini-1.5-Pro 46% 65% 42% 39% 34.95 16.95
Claude-3.5-Sonnet 64% 70% 51% 47% 18.05 10.40
CodeQwen1.5 24% 65% 43% 40% 25.40 6.80
DeepSeek-Coder 37% 70% 39% 35% 7.20 2.95
CodeLlama 16% 60% 41% 37% 19.30 3.95
CodeGemma 13% 50% 31% 28% 15.00 2.30
Java GPT-4-Turbo 21% 35% 15% 12% 7.05 2.20
GPT-3.5-Turbo 13% 25% 8% 7% 7.50 0.80
GPT-o1 41% 60% 44% 35% 15.70 6.85
Gemini-1.5-Pro 19% 30% 14% 12% 23.30 3.90
Claude-3.5-Sonnet 53% 75% 47% 33% 12.35 7.30
CodeQwen1.5 0% 0% 0% 0% 12.95 0.00
DeepSeek-Coder 8% 20% 5% 5% 7.00 0.60
CodeLlama 0% 0% 0% 0% 7.85 0.00
CodeGemma 0% 0% 0% 0% 10.50 0.00
JavaScript GPT-4-Turbo 67% 75% 56% 46% 16.30 11.10
GPT-3.5-Turbo 51% 65% 37% 28% 13.25 8.05
GPT-o1 87% 95% 87% 75% 39.40 33.30
Gemini-1.5-Pro 59% 70% 64% 61% 45.85 22.55
Claude-3.5-Sonnet 65% 80% 59% 53% 20.25 13.35
CodeQwen1.5 23% 35% 25% 20% 8.45 4.80
DeepSeek-Coder 62% 85% 50% 35% 11.85 7.90
CodeLlama 26% 85% 20% 14% 48.75 18.00
CodeGemma 29% 55% 28% 21% 9.00 3.00
The main results of the LLMs’ unit test generation performance focus on the vanilla unit tests extracted directly from the LLMs’ outputs without any changes. The goal of this scenario is to assess the LLMs’ current raw capability to generate project-level unit tests.
Table 2 shows the evaluation results of the vanilla unit tests. First, we observe that different LLMs have varying language-level expertise. For example, Claude-3.5-Sonnet performs the best in Java but falls behind GPT-o1 on JavaScript. Second, we can see from the results that LLMs having different metric-level expertise as well, validating the effectiveness of different evaluation metrics. For example, in Python, Claude-3.5-Sonnet performs the best on CR and ComR while falling behind GPT-o1 on LC and BC.
Among three programming languages, Java is the most difficult language, primarily due to stricter syntax. Many models fail to generate valid Java code, leading to low compilation rates and execution coverage. Among all the evaluated models, GPT-o1 performs the best in general, especially in JavaScript. CodeLlama and CodeGemma have the worst general performance. We also observe that some models tend to generate more unit tests. However, generating more unit tests does do not necessarily lead to better coverage rates. For example, Gemini-1.5-Pro tends to generate the most unit tests but does not obtain the best coverage rate. Additionally, we observe that sometimes the open-source model can even outperform some closed-source models. For example, DeepSeek-Coder works better than GPT-3.5-Turbo on Python and JavaScript. Finally, we confirmed from such results that dependencies exist in metrics. On Java, models like CodeQwen1.5, CodeLlama, and CodeGemma fail to generate compilable unit tests, resulting in the lowest correctness rates and coverage rates.
5.2 Manual Fixing Results
Table 3: Manual Fixing Results with Improvements. CR represents Correctness Rate; ComR represents Compilation Rate; LC represents Line Coverage; BC represents Branch Coverage. The improvements are shown in parentheses.
Language Model CR ComR LC BC #Tests #Correct
Python GPT-4-Turbo 74% (+27%) 100% 65% (+25%) 59% (+23%) 12.60 9.30
GPT-3.5-Turbo 64% (+27%) 100% 63% (+25%) 57% (+23%) 16.90 10.50
GPT-o1 89% (+29%) 100% 88% (+32%) 86% (+32%) 36.35 32.25
Gemini-1.5-Pro 61% (+15%) 100% 71% (+29%) 68% (+29%) 34.95 22.10
Claude-3.5-Sonnet 92% (+28%) 100% 74% (+23%) 70% (+23%) 18.05 16.40
CodeQwen1.5 46% (+22%) 100% 70% (+27%) 65% (+25%) 25.40 10.90
DeepSeek-Coder 53% (+16%) 100% 60% (+21%) 54% (+19%) 7.20 4.10
CodeLlama 31% (+15%) 100% 61% (+20%) 56% (+19%) 19.30 7.20
CodeGemma 36% (+23%) 100% 54% (+23%) 49% (+21%) 15.00 7.85
Java GPT-4-Turbo 59% (+38%) 100% 40% (+25%) 32% (+20%) 7.05 5.05
GPT-3.5-Turbo 54% (+41%) 100% 36% (+28%) 27% (+20%) 7.50 4.55
GPT-o1 64% (+23%) 100% 65% (+21%) 56% (+21%) 15.7 10.75
Gemini-1.5-Pro 56% (+37%) 100% 54% (+40%) 53% (+41%) 23.30 15.25
Claude-3.5-Sonnet 74% (+21%) 100% 60% (+13%) 53% (+20%) 12.35 9.65
CodeQwen1.5 60% (+60%) 100% 42% (+42%) 31% (+31%) 12.95 8.40
DeepSeek-Coder 52% (+44%) 100% 33% (+28%) 19% (+14%) 7.00 3.80
CodeLlama 36% (+36%) 100% 25% (+25%) 20% (+20%) 7.85 4.95
CodeGemma 57% (+57%) 100% 37% (+37%) 22% (+22%) 10.50 6.50
JavaScript GPT-4-Turbo 89% (+22%) 100% 75% (+19%) 59% (+13%) 16.30 14.20
GPT-3.5-Turbo 74% (+23%) 100% 58% (+21%) 45% (+17%) 13.25 11.20
GPT-o1 91% (+4%) 100% 92% (+5%) 79% (+4%) 39.40 35.15
Gemini-1.5-Pro 76% (+17%) 100% 88% (+24%) 80% (+19%) 45.85 33.45
Claude-3.5-Sonnet 87% (+22%) 100% 77% (+18%) 68% (+15%) 20.25 17.55
CodeQwen1.5 32% (+9%) 100% 35% (+10%) 27% (+7%) 8.45 6.15
DeepSeek-Coder 67% (+5%) 100% 58% (+8%) 43% (+8%) 11.85 8.10
CodeLlama 62% (+36%) 100% 44% (+24%) 28% (+14%) 48.75 31.50
CodeGemma 58% (+29%) 100% 50% (+22%) 38% (+17%) 9.00 6.40
Table 3 shows the evaluation results with improvements compared to vanilla results after manual fixing. After fixing compilation and cascade errors, the results show significant improvements across all programming languages and LLMs compared to vanilla unit tests. This indicates that the unit tests generated by LLMs are highly sensitive to compilation and cascade errors.
Among all programming languages, Java benefits the most from manual fixing. In the case of vanilla unit tests, Java exhibits the lowest compilation rates, making it particularly challenging. However, after manual fixing, Java shows the most substantial improvement, highlighting the potential of LLMs for Java after fixing compilation and cascade errors. Among all models, GPT-o1 still performs the best after manual fixing, and CodeLlama and CodeGemma still exhibit the worst general performance. Gemini-1.5-Pro shows the best coverage improvement overall, indicating its strong potential for better unit test generation once compilation and cascade errors are corrected. Finally, we observe that certain models that initially underperform can match or surpass stronger models after manual fixing. For example, in Java, CodeQwen1.5 outperforms DeepSeek-Coder and is now on par with GPT-4-Turbo. In Python, Gemini-1.5-Pro surpasses CodeQwen1.5-7B-Chat, showing better potential after manual fixing. On JavaScript, GPT-3.5-Turbo has reached parity with DeepSeek-Coder.
5.3 LLMs Self-fixing Results
Table 4: Evaluation Results after Self-fixing. CR represents Correctness Rate; ComR represents Compilation Rate; LC represents Line Coverage; BC represents Branch Coverage. The comparisons with manual fixing are shown in parentheses.
Language Model CR ComR LC BC #Tests #Correct
Python GPT-4-Turbo 52% (-22%) 70% (-30%) 39% (-26%) 35% (-24%) 8.85 4.55
GPT-3.5-Turbo 52% (-12%) 75% (-25%) 45% (-18%) 39% (-18%) 14.15 8.20
GPT-o1 67% (-22%) 70% (-30%) 60% (-28%) 58% (-28%) 35.50 24.35
Gemini-1.5-Pro 47% (-14%) 60% (-40%) 45% (-26%) 42% (-26%) 34.95 17.40
Claude-3.5-Sonnet 86% (-6%) 90% (-10%) 67% (-7%) 63% (-7%) 18.00 15.55
CodeQwen1.5 22% (-24%) 60% (-40%) 41% (-29%) 37% (-28%) 25.15 6.25
DeepSeek-Coder 18% (-35%) 35% (-65%) 20% (-40%) 18% (-36%) 4.30 1.45
CodeLlama 0% (-31%) 5% (-95%) 5% (-56%) 5% (-51%) 3.90 0.00
CodeGemma 8% (-28%) 25% (-75%) 14% (-40%) 13% (-36%) 9.15 0.70
Java GPT-4-Turbo 43% (-16%) 55% (-45%) 26% (-14%) 18% (-14%) 6.40 2.80
GPT-3.5-Turbo 17% (-37%) 25% (-75%) 11% (-25%) 12% (-15%) 6.90 1.05
GPT-o1 68% (+4%) 85% (-15%) 58% (-7%) 54% (-2%) 15.60 10.10
Gemini-1.5-Pro 31% (-25%) 40% (-60%) 29% (-25%) 24% (-29%) 22.65 7.15
Claude-3.5-Sonnet 55% (-19%) 70% (-30%) 39% (-21%) 31% (-22%) 10.95 6.70
CodeQwen1.5 5% (-55%) 5% (-95%) 0% (-42%) 0% (-31%) 12.60 0.05
DeepSeek-Coder 13% (-39%) 20% (-80%) 5% (-28%) 2% (-17%) 1.35 0.25
CodeLlama 0% (-36%) 0% (-100%) 0% (-25%) 0% (-20%) 1.30 0.00
CodeGemma 2% (-55%) 5% (-95%) 3% (-34%) 0% (-22%) 1.75 0.05
JavaScript GPT-4-Turbo 70% (-19%) 85% (-15%) 48% (-27%) 35% (-24%) 8.35 6.35
GPT-3.5-Turbo 64% (-10%) 75% (-25%) 40% (-18%) 30% (-15%) 9.70 5.00
GPT-o1 54% (-37%) 65% (-35%) 47% (-45%) 38% (-41%) 20.30 12.25
Gemini-1.5-Pro 75% (-1%) 85% (-15%) 71% (-17%) 65% (-15%) 40.95 28.65
Claude-3.5-Sonnet 74% (-13%) 80% (-20%) 60% (-17%) 53% (-15%) 18.05 13.35
CodeQwen1.5 55% (+23%) 95% (-5%) 66% (+31%) 52% (+25%) 26.10 15.50
DeepSeek-Coder 14% (-53%) 35% (-65%) 15% (-43%) 10% (-33%) 2.90 1.00
CodeLlama 9% (-53%) 35% (-65%) 7% (-37%) 5% (-23%) 7.15 0.55
CodeGemma 31% (-27%) 60% (-40%) 29% (-21%) 21% (-17%) 10.85 3.05
During LLM self-fixing, conversation history and error messages are provided to help the model correct errors. This scenario assesses the LLM’s ability to fix its own mistakes and its potential to generate better unit tests by incorporating self-fixing.
Table 4 shows the LLM self-fixing evaluation results in comparison with manual fixing. First, we observe that most closed-source models have the ability to self-fix errors and generate better unit tests compared to vanilla results, while the evaluated open-source models lack reliable self-fixing abilities. This limitation may stem from factors such as restricted input length, which leads to incomplete context, as well as weaker comprehension and instruction-following abilities. For instance, open-source models like CodeGemma and CodeLlama tend to generate textual instructions for fixing errors rather than directly producing the corrected unit tests specified in the prompt.
Second, we observe that LLM self-fixing follows similar but not identical trends to manual fixing, suggesting that while LLMs’ potential for improvement generally aligns with self-fixing capabilities, some LLMs do not follow this pattern. For instance, in JavaScript, GPT-o1’s self-fixing performance on the coverage rate is significantly worse compared to manual fixing due to a lower number of generated unit tests and a reduced compilation rate.
Although LLM self-fixing currently lags behind manual fixing in performance, LLM self-fixing still holds significant potential. Self-fixing has proven effective when LLMs have the necessary capabilities, and it even has the potential to surpass manual fixing due to its flexibility. For example, in JavaScript, CodeQwen1.5 shows greater improvement through self-fixing compared to manual fixing. This is because, in its vanilla output, CodeQwen1.5 sometimes fails to understand the prompt and does not generate any unit tests. Manual fixing based solely on these initial outputs cannot resolve this issue. However, LLM self-fixing can overcome this limitation by correctly interpreting unit test generation prompts when error messages indicate that unit tests are required.
5.4 Unique Contribution of Unit Tests
We also explore the unique contribution of the generated unit tests on Python. The unique contribution is defined as the total portion of coverage contributed by each generated unit test that does not overlap with the coverage of other unit tests. This is important for several reasons. First, some LLMs generate more unit tests than others, making it insufficient to rely solely on coverage rate as a metric; the unique contribution of each test should also be considered. Second, it’s crucial for LLMs to generate fewer unit tests while still achieving a high coverage rate, as running a large number of tests can sometimes be resource- or time-intensive.
Table 5 reveals that all the tested LLMs have low rates of unique contributions, indicating a tendency to produce redundant and repetitive unit tests. Although GPT-o1 has better coverage rates than Claude-3.5-Sonnet, GPT-o1 produces significantly more unit tests, and its unique contribution is lower than Claude-3.5-Sonnet’s, indicating that it relies on quantity rather than quality to reach higher coverage. As a result, this approach may compromise the overall efficiency of the testing process.
Table 5: Unique Contribution on Vanilla Unit Tests.
Model #Tests LC BC Unique Contribution
GPT-4-Turbo 12.60 40% 36% 6.35%
GPT-3.5-Turbo 16.90 38% 34% 5.90%
GPT-o1 36.35 56% 54% 6.75%
Gemini-1.5-Pro 34.95 42% 39% 6.05%
Claude-3.5-Sonnet 18.05 51% 47% 11.40%
CodeQwen1.5 25.40 43% 40% 3.75%
DeepSeek-Coder 7.20 39% 35% 8.90%
CodeLlama 19.30 41% 37% 5.55%
CodeGemma 15.00 31% 28% 2.70%
5.5 Error Analyses
We conduct complex analyses of compilation, cascade, and post-fix errors per programming language, highlighting the common errors and potential reasons behind the errors. The full analyses are presented in Appendix E.
Compilation Error Analyses. In Python, common compilation errors arise from incorrect import paths for project functions or classes, hallucinated import names or paths, and mismatched parentheses. Java, a syntax-heavy programming language compared to Python and JavaScript, encounters various compilation errors, like hallucinated methods, constructors, or classes, missing essential elements like package declarations, illegal access to private or protected elements, invalid code generation, and improper use of mocking frameworks like Mockito, along with argument type mismatches, ambiguous references, and incompatible types. In JavaScript, common errors include hallucinated imports with incorrect paths, empty test suites, and syntax errors from incomplete code generation or mismatched parentheses.
Cascade Error Analyses. For Python, cascade errors include missing imports (e.g., numpy, unittest, project functions/classes) and FileNotFoundError due to unmocked external files. For Java, the most common cascade error is improper or missing mocking of user interactions, leading to unusable coverage reports when tests terminate abruptly. For JavaScript, the cascade errors include missing imports (e.g., chai, three, project functions/classes), confusion between named and default imports, and Jest framework compliance issues.
Post-Fix Error Analyses. For all programming languages, the mismatch between expected and actual values is the most common error. In Python, AttributeError often occurs due to LLMs hallucinating non-existent attributes. In Java, frequent errors include NullPointerException, zero interactions with mocks, and failures to release mocks due to improper usage. Another frequent error in JavaScript is TypeError, typically caused by LLMs hallucinating non-existent functions and constructors or LLMs invalidly mocking some variables.
Overall. Common errors across different programming languages include hallucinations of functions or classes, and missing required functions or classes. Missing required functions or classes often occurs because LLMs prioritize logical structure over boilerplate code and fail to understand the codebase structure and the dependencies between functions, classes, or modules. Failure to understand the codebase structure and dependencies can also cause other mistakes, such as confusing non-package and package-based projects (Python) or incorrectly using functions, classes, or packages (Java). The most common post-fix error is the mismatch between expected and received values, often caused by incorrect expected values due to the weak reasoning abilities of LLMs.
6 Conclusion
In conclusion, we build a reliable and high-quality project-level unit test generation benchmark – ProjectTest – with three programming languages. We comprehensively evaluate nine LLMs’ unit test generation abilities with/without manual fixing and LLM self-fixing mechanism on ProjectTest. Besides, we conduct comprehensive error analyses per programming language.
Limitations
Our study has several limitations. First, due to our capacity, we mainly focus on three programming languages—Python, Java, and JavaScript—missing the chance to include other languages like C and C#. Additionally, given the fact that the input length restrictions of current LLMs make them unsuitable for handling larger projects in their entirety, we selected moderate-sized projects, allowing us to explore issues like the robustness of LLMs in unit test generation (e.g., hallucinations or incorrect assertions) rather than focusing solely on their ability to handle long-context inputs.
Appendix A Dataset
We provide the detailed information of our datasets in Table 6, Table 7, and Table 8. We provide programming language, project name, license, link, number of stars, and number of forks for each individual project.
Table 6: Dataset Details (Python).
Project Name License Link #Stars #Forks
blackjack MIT license blackjack 2937 641
bridge MIT license bridge 2937 641
doudizhu MIT license doudizhu 2937 641
fuzzywuzzy MIT license fuzzywuzzy 9200 876
gin_rummy GPL-2.0 license gin_rummy 2937 641
keras_preprocessing MIT license keras_preprocessing 1024 443
leducholde MIT license leducholde 2937 641
limitholdem MIT license limitholdem 2937 641
mahjong MIT license mahjong 2937 641
nolimitholdem MIT license nolimitholdem 2937 641
slugify MIT license slugify 1500 109
stock CC-BY-SA-4.0 license stock 10700 1800
stock2 CC-BY-SA-4.0 license stock2 10700 1800
stock3 CC-BY-SA-4.0 license stock3 10700 1800
stock4 CC-BY-SA-4.0 license stock4 10700 1800
structly CC-BY-SA-4.0 license structly 10700 1800
svm MIT license svm 10800 1800
the fuzz CC-BY-SA-4.0 license the fuzz 2949 141
tree CC-BY-SA-4.0 license tree 10800 1800
uno MIT license uno 2937 641
Table 7: Dataset Details (Java).
Project Name License Link #Stars #Forks
Actor_relationship_game Apache-2.0 license Actor_relationship_game 85 5
banking application MIT license banking application 341 366
CalculatorOOPS MIT license CalculatorOOPS 525 513
emailgenerator MIT license emailgenerator 525 513
heap MIT license heap 60500 19600
idcenter Apache-2.0 license idcenter 146 136
libraryApp MIT license libraryApp 341 366
libraryManagement MIT license libraryManagement 341 366
logrequestresponseundertow Author Permission logrequestresponseundertow 152 131
Password_Generator MIT license Password_Generator 341 366
Pong Game MIT license Pong Game 341 366
redis Apache-2.0 license redis 413 218
servlet MIT license servlet 341 366
simpleChat MIT license simpleChat 543 1500
springdatamongowithcluster Author Permission springdatamongowithcluster 152 131
springmicrometerundertow Author Permission springmicrometerundertow 152 131
springreactivenonreactive Author Permission springreactivenonreactive 152 131
springuploads3 Author Permission springuploads3 152 131
Train MIT license Train 545 1600
The license of "Author Permission" in Table 7 means that we obtain the usage permission from the author of the corresponding repository [^7] https://github.com/frandorado/spring-projects/tree/master.
Table 8: Dataset Details (JavaScript).
Project Name License Link #Stars #Forks
aggregate MIT license aggregate 1500 18
animation MIT license animation 103000 35400
check MIT license check 1500 18
circle MIT license circle 2700 330
ckmeans ISC license ckmeans 3400 226
controls MIT license controls 103000 35400
convex MIT license convex 2700 330
easing MIT license easing 418 9
magnetic MIT license magnetic 418 9
overlapkeeper MIT license overlapkeeper 2700 330
particle MIT license particle 2700 330
pixelrender MIT license pixelrender 2400 274
plane MIT license plane 2700 330
solver MIT license solver 2700 330
span MIT license span 2400 274
spherical MIT license spherical 103000 35400
synergy MIT license synergy 310 3
t_test ISC license t_test 3400 226
validate MIT license validate 1500 18
zone MIT license zone 2400 274
Appendix B More Implementation Details
B.1 Prompts
{mdframed}
[style=prompt, backgroundcolor=white] System Prompt: You are a coding assistant. You generate only source code.
User Prompt: {Original Codes} Please generate enough unit test cases for each java file in the {method_signature} project. Ensure to use mock properly for unit tests. Make sure the tests can successfully compile. Make sure the tests have correct results. Try to achieve the highest coverage rate.
Figure 7: The prompt used to generate unit tests for Java projects.
{mdframed}
[style=prompt, backgroundcolor=white] System Prompt: You are a coding assistant. You generate only source code.
User Prompt: {Original Codes} Please generate enough unit test cases for every javascript file in {method_signature} project. Make sure the tests can successfully compile. Make sure the tests have correct results. Try to achieve the highest coverage rate.
Figure 8: The prompt used to generate unit tests for JavaScript projects.
{mdframed}
[style=prompt, backgroundcolor=white] System Prompt: You are a coding assistant. You generate only source code.
[…]
Figure 9: The prompt used to generate unit tests for Python projects.
The prompts are displayed in Figure 7,  8, and  9.
B.2 Models
Table 9: Model Details.
Model Type Model Name License Link
Close-sourced GPT-4-Turbo - https://platform.openai.com/docs/models/gpt-4#gpt-4-turbo-and-gpt-4
Close-sourced GPT-3.5-Turbo - https://platform.openai.com/docs/models/gpt-4#gpt-3-5-turbo
Close-sourced GPT-o1 - https://platform.openai.com/docs/models#o1
Close-sourced Gemini-1.5-Pro - https://ai.google.dev/gemini-api/docs/models/gemini#gemini-1.5-pro
Close-sourced Claude-3.5-Sonnet - https://www.anthropic.com/claude/sonnet
Open-sourced CodeQwen1.5-7B-Chat Tongyi Qianwen LICENSE AGREEMENT https://huggingface.co/Qwen/CodeQwen1.5-7B-Chat
Open-sourced DeepSeek-Coder-6.7b-Instruct DEEPSEEK LICENSE AGREEMENT https://huggingface.co/deepseek-ai/deepseek-coder-6.7b-instruct
Open-sourced CodeLlama-7b-Instruct-hf LLAMA 2 COMMUNITY LICENSE AGREEMENT https://huggingface.co/codellama/CodeLlama-7b-Instruct-hf
Open-sourced CodeGemma-7b-it Gemma Terms of Use https://huggingface.co/google/codegemma-7b-it
The detailed information of models, including license and link, is provided in Table 9.
Appendix C More Statistics
Table 10: Percentages of the vanilla unit tests containing expected and actual value comparisons.
Model GPT-4-Turbo GPT-3.5-Turbo GPT-o1 Gemini Claude CodeQwen DeepSeek-Coder CodeLlama CodeGemma
Python 98% 99% 98% 89% 99% 97% 96% 99% 88%
Java 97% 90% 98% 98% 97% 89% 94% 85% 93%
JavaScript 100% 89% 96% 100% 100% 100% 96% 86% 100%
Table 10 presents the percentages of the vanilla-generated unit tests containing comparisons between expected and actual values per language and per model.
Appendix D Ablation Study
D.1 Ablation Study on Prompts
Table 11: Ablation Study. The Performance of Unit Test Generation by GPT-4-Turbo Using Different Prompts.
Phase Settings CR ComR LC BC #Tests #Correct Tests
Vanilla Full Prompt 47% 65% 40% 36% 12.60 6.15
w/o CR 33% $\downarrow$ 65% 42% 38% 12.75 4.75
w/o ComR 35% 63% $\downarrow$ 41% 38% 11.20 3.95
w/o Coverage 43% 75% 46% $\uparrow$ 42% $\uparrow$ 9.80 4.20
w/o PL 47% 75% 53% 49% 9.95 4.35
w/ Comments 41% 65% 45% 41% 10.65 4.15
Manual Fixing Full Prompt 74% 100% 65% 59% 12.60 9.30
w/o CR 76% $\uparrow$ 100% 69% 64% 12.75 9.90
w/o ComR 75% 100% 70% 65% 11.20 8.35
w/o Coverage 68% 100% 66% $\uparrow$ 61% $\uparrow$ 9.80 6.75
w/o PL 70% 100% 70% 66% 9.95 6.90
w/ Comments 66% 100% 68% 62% 10.65 7.00
We perform a detailed ablation study to analyze the impact of prompts on the performance of unit test generation by LLMs. As mentioned in § 3.3, the prompt is composed of programming language-specific requirements (PL), as well as requirements related to the correctness rate (CR), the compilation rate (ComR), and the coverage rate metrics (Coverage). We ablate each component and analyze the performance of unit test generation of GPT-4-Turbo using different prompts as shown in Table 11. Requirements related to CR and ComR can help improve performance in vanilla unit tests. Coverage-related requirements are not always beneficial, possibly because a high coverage rate is too abstract for LLMs to interpret effectively. Programming language-specific requirements improve performance in CR but have the opposite effect on ComR, LC, and BC.
Besides, we follow the prompt template from previous work like Siddiq et al. (2024) to move the prompts into comments (e.g., /*…*/). We compare the performance with and without comment signs in Table 11. Experimental results show that our prompt demonstrates a significant advantage in CR, while the prompt with comment signs exhibits marginal advantages in ComR, LC, and BC.
D.2 Effect of Compilation Errors and Cascade Errors
Table 12: Evaluation Results When Only Manually Fixing Compilation Errors
Language Model CR ComR LC BC #Tests #Correct Tests
Python GPT-4-Turbo 73% 100% 65% 59% 12.60 9.10
GPT-3.5-Turbo 63% 100% 62% 56% 16.90 10.40
GPT-o1 89% 100% 88% 85% 36.35 32.25
Gemini-1.5-Pro 61% 100% 71% 68% 34.95 22.10
Claude-3.5-Sonnet 92% 100% 74% 70% 18.05 16.40
CodeQwen1.5 40% 100% 65% 59% 25.40 9.60
DeepSeek-Coder 53% 100% 60% 54% 7.20 4.10
CodeLlama 26% 100% 56% 50% 19.30 6.15
CodeGemma 30% 100% 52% 47% 15.00 6.15
Java GPT-4-Turbo 59% 100% 42% 34% 7.05 5.05
GPT-3.5-Turbo 48% 100% 37% 29% 7.50 4.20
GPT-o1 62% 100% 67% 56% 15.70 10.50
Gemini-1.5-Pro 55% 100% 54% 53% 23.30 15.00
Claude-3.5-Sonnet 73% 100% 63% 57% 12.35 9.60
CodeQwen1.5 49% 100% 49% 39% 12.95 7.50
DeepSeek-Coder 40% 100% 36% 19% 7.00 2.85
CodeLlama 30% 100% 26% 21% 7.85 4.25
CodeGemma 46% 100% 44% 26% 10.50 5.55
JavaScript GPT-4-Turbo 89% 100% 75% 59% 16.30 14.15
GPT-3.5-Turbo 71% 100% 56% 44% 13.25 10.65
GPT-o1 91% 100% 92% 79% 39.40 35.15
Gemini-1.5-Pro 76% 100% 88% 80% 45.85 33.30
Claude-3.5-Sonnet 83% 100% 75% 66% 20.25 16.75
CodeQwen1.5 28% 100% 29% 22% 8.45 5.65
DeepSeek-Coder 66% 100% 58% 43% 11.85 8.05
CodeLlama 28% 100% 20% 15% 48.75 21.40
CodeGemma 45% 100% 43% 30% 9.00 5.75
We manually fix only compilation errors and evaluate the corrected unit tests in Table 12.
By fixing compilation errors, Table 12 shows significant improvements across all programming languages and LLMs compared to Table 2, indicating that all the programming languages and LLMs are highly sensitive to compilation errors. Comparing Table 12 with Table 3, we can observe that CodeQwen1.5, CodeGemma, and CodeLlama are more sensitive to cascade errors. For Java, the changes in Table 3 compared to Table 12 are primarily due to missing or invalid mocks of user interactions [^8] We consider coverage rates as not applicable when requiring user interactions. which occur more frequently in unit tests generated by CodeQwen1.5 and CodeGemma.
Appendix E Detailed Error Analyses
We conduct complex analyses of compilation, cascade, and post-fix errors, highlighting the common errors and potential reasons behind the errors.
Compilation Error Analyses
Figure 10 highlights the detailed compilation errors that occurred. One of the most common compilation errors in Python arises from the LLM’s inability to determine whether the project being tested is a package. Specifically, LLMs struggle to recognize the presence or absence of __init__.py files, which define a package, leading to confusion between package-based and non-package projects. This inability leads LLM to fail to correctly import functions or classes from the tested project. Other compilation errors include hallucinating the paths or names of imported functions/classes and mismatched parentheses. Java, a syntax-heavy programming language compared to Python and JavaScript, encounters various compilation errors, resulting in a significantly lower compilation rate than other languages. Java compilation errors often arise from issues like hallucinated methods, constructors, or classes, such as incorrect or non-existent imports and references. Missing essential information, such as required functions, classes, or packages, and package declarations, is also a common problem. Errors frequently occur due to illegal access to private or protected elements, invalid code generation (e.g., generating text instead of code), and improper use of mocking frameworks like Mockito, including incorrect objects, missing or misused MockMvc injections, and argument mismatches. Other errors include incorrect usage of other functions, classes, or packages—such as argument type errors, ambiguous references, or incompatible types. One of the most common compilation errors in JavaScript is the hallucination of imported functions or classes, where the issue often lies in incorrect paths for the imported functions or classes. CodeQwen1.5 has a particularly common compilation error involving invalid generation. This typically occurs due to difficulty understanding the prompt, the need for more specific or detailed code requirements, or the assumption that the code is part of a larger project, leading it to decline generating unit tests. Other compilation errors include test suites containing empty unit tests and syntax errors caused by incomplete code generation or mismatched parentheses.
Cascade Error Analyses
Figure 11 highlights the detailed cascade errors that occurred. For Python, the cascade errors include missing imports of commonly used packages such as numpy and unittest, missing imports of functions or classes from the tested project, and FileNotFoundError. For Java, the most common cascade error is missing or invalid mocking of user interactions. A proper unit test should simulate user interactions through mocking rather than relying on real user inputs. This issue also results in unusable coverage reports for some tested projects, as the error forces an abrupt termination, preventing the generation of coverage data. For JavaScript, the cascade errors include missing imports of commonly used packages such as chai and three, and missing imports of functions or classes from the tested project. Two other common errors specific to JavaScript are that LLMs may confuse named imports with default imports and fail to comply with the Jest framework.
Post-Fix Error Analyses
Figure 12 highlights the incorrectness reasons after all manual fixes. For all programming languages, the mismatch between expected and actual values (AssertionError) is the most common error. Another frequent error in Python is AttributeError, typically caused by LLMs hallucinating non-existent attributes. Other frequent problems in Java include NullPointer Errors, zero interactions with mocks, and failures to release mocks, often due to improper mock usage. For projects tested with the Spring framework, errors specific to Spring are also common. Another frequent error in JavaScript is TypeError, mostly caused by LLMs hallucinating non-existent functions and constructors or LLMs invalidly mocking some variables.
{forest}
forked edges, for tree= grow’=0, draw, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=4.4em,font=,, where level=2text width=12em,font=,, where level=3text width=25em,font=,, [ Compilation Error Analysis, ver [ Python, fill=lgreen [ Confuse between non-package and package-based projects , leaf, text width=28em, fill=lgreen ] [ Hallucinate the imported functions/classes:
1. Paths of the imported functions/classes are wrong
2. Names of the imported functions/classes are wrong , leaf, text width=28em, fill=lgreen ] [ Syntax Error: Mismatched parentheses , leaf, text width=28em, fill=lgreen ] ] [ Java, fill=lblue [ Hallucinate methods/constructors/functions/classes:
1. Paths of the imported functions/classes are wrong
2. Names of the imported functions/classes are wrong
3. Non-existed methods/constructors , leaf, text width=28em, fill=lblue ] [ Missing information:
1. Required functions/classes/packages are missing
2. Required package information is missing
3. Unreported exception
, leaf, text width=28em, fill=lblue ] [ Illegal access to private/protected functions/classes , leaf, text width=28em, fill=lblue ] [ Invalid generation:
1. Generate textual instructions instead of codes 2. Block by model
, leaf, text width=28em, fill=lblue ] [ Incorrect use of mocking:
1. Wrong objects provided to Mockito
2. Missing MockMvc injection 3. Inappropriate mockmvc
4. Argument mismatch , leaf, text width=28em, fill=lblue ] [ Incorrect use of other functions/classes/packages:
1. Arguments type error 2. Ambiguous reference
3. Incompatible types , leaf, text width=28em, fill=lblue ] ] [ JavaScript, fill=lyellow[ Hallucinate the imported functions/classes:
1. Paths of the imported functions/classes are wrong , leaf, text width=28em, fill=lyellow ] [ Invalid generation:
1. Cannot understand the prompt 2. Require more/specific codes
3. Assume the codes are part of a larger project and
decline to generate unit tests , leaf, text width=28em, fill=lyellow ] [ Test suits have empty unit tests
, leaf, text width=28em, fill=lyellow ] [ Syntax Error:
1. Incomplete generation 2. Mismatched parentheses , leaf, text width=28em, fill=lyellow ] ] ]
Figure 10: Frequent Compilation Errors in Main Results.
{forest}
forked edges, for tree= grow’=0, draw, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=4.4em,font=,, where level=2text width=12em,font=,, where level=3text width=20em,font=,, [ Cascade Error Analysis, ver [ Python, fill=lgreen [ Required functions/classes/libraries are missing:
1. Import numpy or unittest.mock
2. Import functions/classes of the tested project , leaf, text width=25em, fill=lgreen ] [ FileNotFoundError , leaf, text width=25em, fill=lgreen ] ] [ Java, fill=lblue [ Missing/Invalid mock of user interactions , leaf, text width=25em, fill=lblue ] ] [ JavaScript, fill=lyellow [ Required functions/classes/libraries are missing:
1. Import chai or three
2. Import functions/classes of the tested project , leaf, text width=25em, fill=lyellow ] [ Confuse between name import and default import , leaf, text width=25em, fill=lyellow ] [ Do not follow the Jest framework , leaf, text width=25em, fill=lyellow ] ] ]
Figure 11: Frequent Cascade Errors.
{forest}
forked edges, for tree= grow’=0, draw, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=4.4em,font=,, where level=2text width=12em,font=,, where level=3text width=25em,font=,, [ Post-fix Error Analysis, ver [ Python, fill=lgreen [ 1. AttributeError 2. AssertionError 3. TypeError 4. ValueError
5. IndexError 6. _csv.Error 7. NameError 8. KeyError 9. Others , leaf, text width=30em, fill=lgreen ] ] [ Java, fill=lblue [ 1. Mismatch between expected and received 2. NullPointer Error
3. Zero interactions with mock 4. Failed to release mocks
5. MissingMethodInvocation 6. Misplaced or misused argument matcher
7. Spring framework error 8. NoSuchElement 9. Others , leaf, text width=30em, fill=lblue ] ] [ JavaScript, fill=lyellow [ 1. Mismatch between expected and received 2. TypeError 3. RangeError
4. RuntimeError 5. ReferenceError 6. SyntaxError 7. Others , leaf, text width=30em, fill=lyellow ] ] ]
Figure 12: Frequent Post-Fix Errors.